### v1

# mode=buildnew

Abstract
When a neural network parameterized loss function consists of many terms, the combinatorial choice of weight multipliers during the optimization process forms a challenging problem. To address this, we proposed a probabilistic graphical model (PGM) for the joint model parameter and multiplier evolution process, with a hypervolume based likelihood that promotes multi-objective descent of each loss term. The corresponding parameter and multiplier estimation as a sequential decision process is then cast into an optimal control problem, where the multi-objective descent goal is dispatched hierarchically into a series of constraint optimization sub-problems. The subproblem constraint automatically adapts itself according to Pareto dominance and serves as the setpoint for the low level multiplier controller to schedule loss landscapes via output feedback of each loss term. Our method is multiplier-free and operates at the timescale of epochs, thus saves tremendous computational resources compared to full training cycle multiplier tuning. We applied it to domain invariant variational auto-encoding with 6 loss terms on the PACS domain generalization task, and observed robust performance across a range of controller hyperparameters, as well as different multiplier initial conditions, outperforming other multiplier scheduling methods. We offered modular implementation of our method, admitting custom definition of many loss terms for applying our multi-objective hierarchical output feedback training scheme to other deep learning fields.
1 Introduction
In many deep learning fields including domain generalization (Gulrajani and Lopez-Paz, 2020), the loss function for neural networks amounts to a summation of multiple terms, with a multiplier weighting each term. As the number of terms increases, the space of choice for those multipliers grows dramatically. In addition, when the validation set is drawn from a different distribution than the target domain, the validation performance does not faithfully reflect the target domain performance and can easily reach and maintain a saturated value during training. All these factors complicate and add difficulties to the process of model selection using conventional hyperparameter tuning (Probst et al., 2019) methods.
Low level optimization algorithms for descending single objective neural network losses (e.g. a weighted summation of many terms), such as ADAM (Kingma and Ba, 2015), have been widely used, although their theoretical behavior was analyzed predominantly with convex optimization theory (Kingma and Ba, 2015; Reddi et al., 2019), which does not hold globally (Milne, 2019). From a control theory point of view, these low level optimization algorithms drive the model parameters iteratively, which forms a dynamic system. Due to lack of comprehensive theoretical underpinnings of the loss function property with respect to model parameters, this dynamic system has unreliable open-loop behavior. When considering multi-objective optimization of the many loss terms, this lack of theoretical understanding becomes more pronounced.
Inspired by feedback control theory (Doyle et al., 2013) and feedback optimization in power systems (Picallo et al., 2020; He et al., 2023), we treat the dynamic system of model parameters driven by the low level optimization algorithm as an uncontrolled plant, with model parameter as state and the many loss terms as output. We use state dependent multiplier induced loss landscapes as control input and designed a hierarchical control structure (Dietterich et al., 1998; Kulkarni et al., 2016; Sun et al., 2020) to drive the system towards a multi-objective descent goal, without modifying the internal dynamic of low level neural network optimization algorithms.
This hierarchical control structure allows for a more refined and flexible approach to high-level optimization of neural networks, by automatically balancing the trade-off between a task-specific empirical loss and various regularization losses, which we detail in  Equation 1. We depicted our idea in the control diagram in Figure 1 where the motivation of choosing multiplier induced landscape scheduling as control input is illustrated in Figure 4.
We demonstrated that our method removes the combination curse of dimensionality of multi-dimensional multipliers, via training domain invariant variational auto-encoding (Ilse et al., 2020) on the PACS dataset. Our method only has a small number of hyperparameters for our controllers while achieving rubust performance when changing them. The baselines include different warmup multiplier scheduling strategies, which still need explicit specification of multiplier ultimate values. Their performance varies drastically with changes in those ultimate values. We attribute the improved performance of our method to the automatic trade-off of different loss terms. Our major contributions are:
  • We proposed a probabilistic graphical model (PGM) for depicting the joint multiplier and model parameter adaptation process with a likelihood function imposed, promoting multi-objective descent of each loss term in structured risk. We proposed to transform the problem of parameter inference of this PGM into a sequential decision process through an optimal control formulation.
  • We proposed to solve the sequential decision process for the multi-objective descent goal via breaking the goal into a series of constraint optimization sub-goals which forms a hierarchical controller that automatically adjusts the value of both the multiplier and the model parameters during the training process. The constraint bound updates itself via Pareto dominance. The multi-objective descent of the empirical risk and multidimensional penalty (regularization) loss in Equation 1 offers additional model selection criteria along with validation set performance. Compared to Bayesian optimization, our controller operates at the timescale of model parameter dynamic epoch instead of a complete training, which is more computationally efficient.
  • When treating the neural network model parameter dynamic system driven by low level optimization algorithm as uncontrolled plant, we proposed to use state dependent multiplier induced loss landscapes as control input (see Figure 4) and offered initial theoretical analysis of the multi-objective descent behavior of our closed loop system . We offered modular implementation of our method  [^1] https://github.com/marrlab/DomainLab/tree/mhof and showed robustness of our method with respect to different multiplier initial conditions and controller hyperparameters for domain generalization on the PACS dataset with domain invariant auto-encoding.
2 Preliminaries
2.1 Structural Risk Minimization
Structural Risk Minimization (SRM) is a principle in learning algorithms that balances training performance and model complexity via optimizing the penalized loss in Equation 1.
$\begin{split}L(\theta,\mu,\mathcal{D}_{tr})&=\ell(\theta,\mathcal{D}_{tr})+\mu^{T}R(\theta,\mathcal{D}_{tr})\\ \mu,R(\cdot)&\in\mathbb{R}_{+}^{d}\end{split}$ (1)
In Equation 1, we consider loss $L$ with training data $\mathcal{D}_{tr}$ and model parameters $\theta$. $\ell(\theta,\mathcal{D}_{tr})$ represents the empirical risk and $R(\theta,\mathcal{D}_{tr})$ represents the regularization term of the loss function, also referred to as the penalty function and thus $L$ the penalized loss. $\mathbb{R}_{+}^{d}$ is the $d$ dimensional positive octant. We refer to the multiplier $\mu$ as the penalty multiplier, as a special case of general hyperparameters (e.g. learning rate) in deep learning.
Notation
In the following sections, for brevity, we omit the implicit dependence of $\ell(\cdot)$ and $R(\cdot)$ on $\mathcal{D}_{tr}$ and $\theta$. We also write $R(\theta,\cdot)$ where we need an explicit dependence on $\theta$ but use $\cdot$ only to represent data $\mathcal{D}_{tr}$ and/or other arguments. We use subscript of $R(\cdot)$ to indicate the component of $R(\cdot)$ corresponding to that subscript, e.g., if $\mu=[\beta,\gamma]$, then $R_{\beta}(\cdot)$ corresponds to the component of $R(\cdot)$ associated with $\beta$. We use $X,Y\sim\mathcal{D}$ to represent a random variable $X$ and its supervision signal $Y$ from distribution $\mathcal{D}$ and we use the same symbol $\mathcal{D}$ to denote a corresponding dataset as well. Let $x$ be an instance of $X$, so is $y$ for $Y$.
We use superscript $k$ in bracket to index the optimization iteration, which by default corresponds to one epoch for our method (it can also correspond to one training cycle in Equation 2 to iteratively achieve Equation 3 for conventional hyperparameter optimization, see Remark 3.1).
Definition 2.1.
When optimizing Equation 1 with multiplier $\mu$ iteratively (e.g. with gradient based methods like ADAM), the model parameter evolution dynamic $\theta^{(k+1)}=f_{\theta}(\mu,\theta^{(k)},\mathcal{D}_{tr})$ brings $\theta^{(k)}$ to its next value $\theta^{(k+1)}$. $\theta^{(k+1)}$ then becomes the initial value for further iterations. For notational simplicity, we replace the explicit dependence of this dynamic on training data $\mathcal{D}_{tr}$, multiplier $\mu$ (potentially varying value for each iteration) and/or even the initial model parameter $\theta$ with dots. This convention results in notations $f_{\theta}(\theta,\cdot)$, $f_{\theta}(\mu,\cdot)$ and $f_{\theta}(\cdot)$. We also write $\theta^{+}=f_{\theta}(\mu,\theta,\cdot)$ and $\theta^{+}=f_{\theta}(\theta,\cdot)$ (potentially varying $\mu$ for each iteration) to represent the model parameter dynamic system without explicitly using $k$.
2.2 Multi-objective optimization
Definition 2.2.
[…]
Definition 2.3.
Starting from $\theta^{(0)}$, under dynamic system $\theta^{+}=f_{\theta}(\theta,\cdot)$, we use $s\mathcal{R}\left(\theta^{(0)},f_{\theta}(\cdot)\right)$ to represent the set of model parameters ($\theta$ points) that can be reached. Accordingly, we use $v\mathcal{R}_{\ell(\cdot),R(\cdot)}\left(\theta^{(0)},f_{\theta}(\cdot)\right)$ to represent the corresponding set of multi-objective function values.
Definition 2.4.
If $\theta_{1}\sim\equiv\theta_{2}$ where $\theta_{1}\in s\mathcal{R}(\theta^{(0)},f_{\theta}(\cdot))$ and $\theta_{2}\in s\mathcal{R}(\theta^{(0)},f_{\theta}(\cdot))$ as in Definition 2.3, we define $\theta_{2}\in\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$ the map from a representing element $\theta_{1}$ to its equivalence class, which contains $\theta_{2}$. We use $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$ to represent the set of $\ell(\cdot),R(\cdot)$ values for each element of $\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$
2.3 Hyperparameter tuning
The multiplier $\mu$ in Equation 1 is a hyperparameter that affects the evolution of the model parameters $\theta$ in the optimization process for $L$ defined in Equation 1, which leads to optimized model parameter $\theta_{\mu,\mathcal{D}_{tr}}$ below, where we omit the potential dependence of $\theta_{\mu,\mathcal{D}_{tr}}$ on its initial condition $\theta^{(0)}$ for notational simplicity.
$\theta_{\mu,\mathcal{D}_{tr}}$ $=\arg\min_{\theta}L(\theta,\mu,\mathcal{D}_{tr}).$ (2)
Hyperparameter optimization aims to adjust the multiplier $\mu$ in alignment with specific performance metrics ${O}$ (e.g. validation set accuracy for a classification task) in Equation 3 evaluated on the validation set $\mathcal{D}_{val}$, which lead to the selected hyperparameter $\mu_{\mathcal{D}_{tr},\mathcal{D}_{val}}$.
$\mu_{\mathcal{D}_{tr},\mathcal{D}_{val}}=\arg\min_{\mu}O(\theta_{\mu,\mathcal{D}_{tr}},\mathcal{D}_{val}).$ (3)
The process of hyperparameter optimization in Equation 3 leads to selected model $\theta_{\mu_{\mathcal{D}_{tr},\mathcal{D}_{val}},\mathcal{D}_{tr}}$, which can be conceptualized as a problem of algorithm configuration (Hutter et al., 2007, 2014; López-Ibáñez et al., 2016; Sun et al., 2020) where the hyperparameter configures a machine learning algorithm.
To solve Equation 3 iteratively, each iteration relies on a complete training cycle to optimize Equation 2. The selection of $\mu$ depends on validation set $\mathcal{D}_{val}$, which can only be drawn from a different distribution compared to the target domain for the domain generalization problem (Section 2.4). Additionally, the performance metric $O$ on the in-domain validation set can reach and maintain a saturated value during training with neural networks of high expressive power.
2.4 Domain generalization
Domain generalization aims at enabling the trained neural network to achieve robust generalization to unseen domains exhibiting distribution shifts (Gulrajani and Lopez-Paz, 2020; Sun et al., 2019b). Many domain generalization methods (Ganin et al., 2016; Li et al., 2018; Levi et al., 2021; Carlucci et al., 2019; Ilse et al., 2020; Sun and Buettner, 2021; Rame et al., 2022) promotes this goal via domain invariant representation by adding domain invariant regularization losses $R(\cdot)$ upon task-specific losses $\ell(\cdot)$ (e.g. classification loss) on the training data in Equation 1.
As an example of $\ell(\cdot)+\mu^{T}R(\cdot)$ loss for domain generalization, we label each component of the loss of Domain Invariant Variational Autoencoding (DIVA) (Ilse et al., 2020) in the following equations, where we use $x$ to denote the input instance (e.g. input image), $y$ to denote the corresponding supervision class label, and $d$ the domain label. DIVA (Ilse et al., 2020) aims to disentangle domain specific information with learned features $z_{d}$ and class specific information with $z_{y}$, and has two corresponding classification losses $\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]$ (classifying the correct class label based on $z_{y}$) and $\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ (classifying the correct domain label based on $z_{d}$).
$L(\theta,\mu,x,y,d)=\ell(\cdot)+\mu^{T}R(\cdot)=$ (4)
$\gamma_{y}\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]+$ (5)
$\mu_{recon}\mathbb{E}_{q_{\phi_{d}}(z_{d}|x),q(z_{x}|x),q(z_{y}|x)}\log p_{\theta_{r}}(x|z_{d},z_{x},z_{y})+$ (6)
$[-\beta_{x}KL(q_{\phi_{x}}(z_{x}|x)||p_{\theta_{x}}(z_{x}))]+$ (7)
$[-\beta_{y}KL(q_{\phi_{y}}(z_{y}|x)||p_{\theta_{y}}(z_{y}|y))]+$ (8)
$[-\beta_{d}KL(q_{\phi_{d}}(z_{d}|x)||p_{\theta_{d}}(z_{d}|d))]+$ (9)
$\gamma_{d}\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ (10)
Here, $KL$ denotes Kullback-Leibler divergence, $\mathbb{E}$ denotes expectation, $p(\cdot)$ stands for the prior distribution or the distribution of the generative model, and $q(\cdot)$ stands for the approximate posterior distribution. For more details, refer to (Ilse et al., 2020). In the above situation, we have $\mu=[\mu_{recon},\beta_{x},\beta_{y},\beta_{d},\gamma_{d}]$ and $\theta=[\phi_{y},\phi_{d},\theta_{r},\theta_{d},\theta_{x},\theta_{y},\phi_{x},\phi_{y},\phi_{d},\omega_{d},\omega_{y}]$ with each entry representing the weight (with bias) of a neural network.
In (Ilse et al., 2020), $\mu_{recon}=1.0$ in Equation 6. Additionally, $\gamma_{y}$ and $\gamma_{d}$ are maintained as constants. $\gamma_{y}$ is corresponding to $\ell(\cdot)=\gamma_{y}\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})])$ in Equation 5, and $\gamma_{d}$ in Equation 10 is associated with one component of $R(\cdot)$. However, the combined choice of $\gamma_{y}$ and $\gamma_{d}$ significantly influences the generalization performance, as shown by our experimental findings in Section 4.
For the multiplier $\beta_{x},\beta_{y},\beta_{d}$ corresponding to other components of $R(\cdot)$, Ilse et al. (2020) used a warm-up strategy to increase them gradually from a small value to a pre-defined value. This, however, still requires a choice of the ultimate values for $\beta_{x},\beta_{y},\beta_{d}$. In (Ilse et al., 2020), these ultimate values are simply set to one. We coin this kind of multiplier scheduling strategy a feedforward scheme. Note that this strategy does not reduce the number of hyperparameters to be selected, since ultimate values for the multipliers still have to be specified.
2.5 Related work
Multipliers warmup (Sønderby et al., 2016; Ilse et al., 2020) or phase-in (Sicilia et al., 2023) requires ultimate multiplier values for each loss term, thus necessitates hyperparameter (ultimate multiplier values) optimization. Other works, such as (Rezende and Viola, 2018; Klushyn et al., 2019), has introduced constraint optimization by representing the weights of loss terms using Lagrange multipliers, designed to prevent over-regularisation of a single term within the loss function, which is extended to multiple loss terms in (Chen et al., 2022). In addition, Shao et al. (2020) (only deals with 2 loss terms),Chen et al. (2022) proposed similar multiplier adaptation schemes as in Section 3.2.3 where the multiplier changes with a rate proportional to how the corresponding loss component distances from a constraint bound. However, their methods relied on a fixed constraint bound  for each loss term and fixed proportionate gain for each loss term (defined to be $K_{I}$ in Equation 48 in our work), necessitating a hyperparameter search, which is challenging when the number of terms in the loss increases. Moreover, for some regularizers, such as the ones based on KL divergence, it is challenging to estimate the constraint bound range. Specifically in the field of domain generalization, it is not obvious how to choose the constraint bound value and other hyperparameters due to the lack of observations from the target domain, where the validation set from the training domain can easily reach a saturated value due to the high expressive power of modern neural networks.
In comparison, we deal with many loss terms with multidimensional multipliers and introduced Equation 23 to progressively adapt the constraint bound and Remark 3.8 to choose the multidimensional controller gain value, thus avoid exhaustive hyperparameter search. In addition to the control theory formulation for multiplier adjustment, we also provided a Probabilistic Graphical Model interpretation in Section 3.1 and a landscape scheduling interpretation in Figure 4. Furthermore, we discussed our method under the multi-objective optimization scheme compared to single-objective consideration from other works.
3 Methods
In this section, we elaborate on the hierarchical output feedback control strategy for the multidimensional multiplier adaptation in Equation 1 (see Figure 1) towards multi-objective optimization. We propose a Probabilistic Graphical Model (PGM) for the joint model parameter and multiplier adaptation process in Section 3.1, which leads to an optimal control formulation of parameter estimation, extending the classical hyperparameter tuning-based approach to parameter-multiplier co-evolution in Section 3.2.
Figure 1: A control diagram illustrating the hierarchical output feedback optimization process we proposed with multi-objective setpoint adaptation. The uncontrolled plant corresponds to an open loop dynamic $\theta^{+}=f_{\theta}(\mu,\theta,\cdot)$ via optimizing Equation 1 with a low level optimization algorithm (e.g. ADAM) lacking feedback. The $\mu$ controller adjusts the multiplier $\mu$, thus schedules loss landscape (see Illustration in Figure 4), based on the difference $e$ between the setpoint $b$ (defined in Equation 23) and the measured output components $R(\cdot)$, guiding the optimization of model parameters $\theta$ through a feedback loop. The setpoint $b$ is adjusted via feedback from $\ell(\cdot)$ and $R(\cdot)$ (Equation 24a) which forms a higher hierarchy.
3.1 Probabilistic graphical model for parameter and multiplier adaptive optimization process
Figure 2: A probabilistic graphical model for the sequential decision process of joint multiplier and model parameter adaptation in multiple-domain (data-site specific distribution) supervised learning.
We propose Figure 2 to describe the sequential decision process of model parameter and multiplier adaptation in the form of a probabilistic graphical model (PGM) (Koller and Friedman, 2009). In the upper part of Figure 2, we model the generative process of $N$ observations of the instance $X$ (e.g. an image) and the corresponding supervision signal $Y$ (e.g. a class label). The path $\beta\rightarrow D_{ob}\rightarrow X$, where, $D_{ob}$ is the data-site Sun et al. (2019a) specific prior parameter (domain variable, characterizing distribution) with hyper-prior $\beta$, models the information which is not captured by the supervision signal $Y$ of $X$.
In the lower part of the graphical model, with superscript $k$ representing the iteration for the decision process, we use $\mu^{(k)}$ to represent the multiplier of the loss in Equation 1 at the $k$th iteration. The multiplier $\mu^{(k)}$ serves as a configuration parameter for the optimization process of Equation 1 that affects the value of $\theta_{\mu}^{(k)}$, evolved from its previous value $\theta^{(k-1)}$. $\theta_{\mu}^{(k)}$ then becomes the initial value for the next iteration. This can be described in Equation 14 with explicit dependence of $\theta^{(k)}$ on $\mu^{(k)}$ omitted. The model parameter $\theta^{(k)}$ and observed data $X$, supervision signal $Y$ co-parent the performance indicator as output ${O}^{(k)}$.
The plate replicates $B$ in the lower part represents the number of optimization iterations (each iteration correspond to $N$ observations). The adaptive generation of the next multiplier $\mu^{(k+1)}$ depends on the previous value $\mu^{(k)}$ (the self-loop dashed arrow in Figure 2) and the feedback information of the output $O^{(k)}$ (long dashed arrow), as in Equation 11.
$\mu^{(k+1)}=f_{\mu}(\mu^{(k)},O^{(k)},\cdot)$ (11)
Remark 3.1.
An example of $f_{\mu}$ can be based on the acquisition function in Bayesian optimization (BO) (Garnett, 2023) for approximating Equation 3, which leads to a BO realization of this PGM in (Sun et al., 2019a), where $O$ in Equation 11 is usually chosen to be the validation set prediction performance and iteration index $k$ corresponds to a whole training cycle of Equation 2 to iteratively achieve Equation 3.
Different from the BO realization of this PGM, which generates multipliers at a much slower timescale (usually once at the end of training) than that for the evolution of model parameters in neural network training, the hierarchical output feedback optimization we proposed (Figure 1 and elaborated in following sections) operates at the timescale of epoch, thus forms another realization of this PGM with iteration index $k$ corresponding to epochs. In addition, we have $O(\cdot)=\ell(\cdot),R(\cdot)$ for the PGM in Figure 2 in contrast to validation set performance used in BO.
We introduce an objective function for estimating $\theta_{\mu^{(B)}}$ based on the profile likelihood in eq. 12 in accordance with the PGM in Figure 2. This likelihood promotes a high $e\mathcal{HV}(\cdot)$ in Definition 3.1 with respect to $\theta_{\mu^{(B)}}$. $\Xi$ is the normalization factor.
$\mathcal{P}(\theta_{\mu^{(B)}})$ $=\frac{1}{\Xi}\exp\left(e\mathcal{HV}_{O(\cdot)}(\theta_{\mu^{(B)}},\theta^{(0)},f_{\theta},\cdot)\right)$ (12)
Definition 3.1.
$e\mathcal{HV}_{O(\cdot)=\ell(\cdot),R(\cdot)}(\theta,\theta^{(0)},f_{\theta},\cdot)$ maps $\theta$ to the dominated hypervolume (Zitzler and Künzli, 2004; Zitzler et al., 2007; Guerreiro et al., 2021) of $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$ in Definition 2.4 with respect to reference point $\ell(\theta^{(0)},\cdot)\in\mathbb{R},R(\theta^{(0)},\cdot)\in\mathbb{R}^{d}$. As illustrated in Figure 3.
Remark 3.2.
The PGM also provides a potential possibility to impose a Bayesian interpretation to the joint multiplier tuning and model training process for frequentist machine learning models, the deeper investigation of which we leave for future work.
Figure 3: Illustration to $e\mathcal{HV}$ in Definition 3.1: We take $[\ell(\theta^{(0)},\cdot)\in\mathbb{R},R(\theta^{(0)})\in\mathbb{R}^{d=1}]$ as reference point. Suppose $\{A,B,C,D,E,F\}$ in the illustration constitute $v\mathcal{R}\left(\theta^{(0)},f_{\theta}(\cdot)\right)$ in Definition 2.3. For any $\theta$ corresponding to the point A in the illustration with coordinate $[\ell(\theta,\cdot),R(\theta,\cdot)]$, $e\mathcal{HV}$ maps $\theta$ to $\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$, then to $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$ (see Definition 2.4), which is the point set {A,B,C,D} , then it calculates the dominated hypervolume with respect to the reference point (union of the shaded rectangles).
3.2 Hierarchical control formulation of joint parameter multiplier optimization
3.2.1 Optimal control formulation
Analogous to how optimal control Lewis et al. (2012) can be formulated as maximum likelihood estimation in a probabilistic graphical model (Levine, 2018; Sun and Bischl, 2019), the estimation of $\theta^{(B)}$ based on the profile likelihood like objective in Equation 12 for Figure 2 is also a sequential decision process. From a statistical inference point of view, we are conducting a sequential inference (i.e., finding a sequence of $\mu$ in Equation 13), such that the last step $B$ has the best possible hyper-volume with respect to output $O$ (in a stochastic sense, see Figure 3), where $O$ can be a performance metric like accuracy on the validation set in Bayesian optimization, or defined in Equation 15 for our method.
To reformulate the estimation problem as an optimal control problem, we present the sequential decision on choosing a sequence of multipliers $\mu^{(1)},\cdots,\mu^{(B)}$ defined in Equation 1 with respect to the objective function in Equation 13 at iteration $B$ which promotes Equation 12.
$\min_{\mu^{(1)},\cdots,\mu^{(k)},\cdots,\mu^{(B)}}$ $e\mathcal{HV}_{O(\cdot)}(\theta_{\mu^{(B)}},\theta^{(0)},f_{\theta},\cdot)$ (13)
$\mathrm{s.t.}\leavevmode\nobreak$ $\theta^{(k)}=f_{\theta}(\mu^{(k)},\theta^{(k-1)},\cdot)$ (14)
$O^{(k)}=\left[\ell(\theta^{(k)}|\cdot),R(\theta^{(k)}|\cdot)\right]$ (15)
$k=0,1,\cdots,B$ (16)
Here, we use $\theta$ to denote the neural network weights, which can be regarded as the state of a dynamic system, where the state transition is governed by a low level model parameter optimization dynamic $f_{\theta}$ in Equation 14 (e.g. ADAM  (Kingma and Ba, 2015)). From a control theory point of view, $f_{\theta}$ depicts the uncertain dynamic of a plant to be controlled, while the task here is to define a controller to generate a $\mu$ sequence to guide the uncertain dynamic of $\theta$ with respect to plant output $O$, as depicted in Figure 1.We treat the loss terms $R(\cdot)$ and $\ell(\cdot)$ as the output of the system in Equation 15. From a statistical point of view, $f_{\theta}$ corresponds to the parent structure of $\theta^{(k)}$ in Figure 2, which drives $\theta$ to the next value.
Although we have no complete knowledge of $f_{\theta}$ from Equation 14 describing low level optimization algorithms like ADAM, in practice, the penalized loss from Equation 1 often descent after some number of iterations (possibly with oscillations in between), which leads to Definition 3.2.
Definition 3.2.
A penalized loss descent operator $\mathcal{G}$ satisfies that if $\theta^{+}\in\mathcal{G}_{\mu}(\theta;\ell(\cdot),R(\cdot))$, then $\ell(\theta^{+})+\mu R(\theta^{+})<\ell(\theta)+\mu R(\theta)$
Therefore, we take the frequent occurrences of operators in Definition 3.2 during neural network training as a mild assumption in our following discussion. For instance, we could run $f_{\theta}$ in Equation 14 in one step to have $\ell(\theta^{(k+1)})+\mu^{(k+1)}R(\theta^{(k+1)})<\ell(\theta^{(k)})+\mu R(\theta^{(k)})$, or in more than one steps to achieve Definition 3.2.
Our final goal, however, is to ensure a joint descent of $\ell(\cdot)$ and $R(\cdot)$, which leads to Definition 3.3:
Definition 3.3.
A Pareto-descent operator takes $\theta$ to $\theta^{(+)}$ (via one or several iterations) which descents $R(\cdot)$ and $\ell(\cdot)$ simultaneously, i.e.,
$\ell(\theta^{+})$ $<\ell(\theta)$ (17)
$R(\theta^{+})$ $\preceq R(\theta)$ (18)
3.2.2 Multi-objective optimization via shrinking the reference signal in constrained optimization
A concrete form of Equation 11 will detail how $\mu$ is updated by a controller $f_{\mu}$. Combining Equations 11 and 14, we have
$\theta^{(k+1)}$ $=f_{\theta}(f_{\mu}(\mu^{(k)},O^{(k)}),\theta^{(k)},\cdot)$ (19)
$=f_{\theta,\mu}^{k}(\mu^{(0)},\theta^{(0)},\cdot)$ (20)
where we use $f_{\theta,\mu}^{k}$ to represent the $k$-times compound recursive evaluations of $f_{\theta}$ and $f_{\mu}$ in Equation 19 until the dependence is only on $\theta^{(0)},\mu^{(0)}$. We call Equation 20 the closed-loop dynamics for $\theta$.
The question then arises: how can we design such a closed-loop dynamic system to ensure a Pareto descent of $O(\cdot)$ (both $R(\cdot)$ and $\ell(\cdot)$), leading to multi-objective optimization in Definition 3.3 or Equation 13?
To resolve this, we reduce it to a simpler problem, i.e., when bounding the regularization term $R(\cdot)$ from above by a time-varying reference $b^{(k)}$ (a.k.a. setpoint) in eq. 26, what value can we achieve for $\ell(\cdot)$ at Equation 25. If we could ensure the bound changes monotonically, multi-objective optimization can be achieved.
First, we define the time-varying reference signal being governed by the following:
$b^{(0)}$ $=\rho R^{(0)}$ (21)
$0<$ $\rho\in\mathbb{R}<1$ (22)
$b^{(k)}$ $=g_{b}(\ell^{(0:k)}(\cdot),R^{(0:k)}(\cdot))$ (23)
where $g_{b}({\cdot})$ represents the mapping from the previous value at the training step $k-1$ to the new value at step $k$, defined as
$b^{(k)}=$ $R^{(k)},\leavevmode\nobreak\ \text{if }R^{(k)}\prec b^{(k-1)}\land\ell^{(k)}<\min_{j}\ell^{(j)}$ (24a)
$b^{(k)}=$ $j=0,\cdots,k-1$ (24b)
$b^{(k)}=$ $b^{(k-1)},\leavevmode\nobreak\ \text{otherwise}$ (24c)
Remark 3.3.
Note that although $b^{(k)}$ is only defined for $R^{(k)}(\cdot)$, the shrinkage of $b^{(k)}$ also depends on $\ell(\cdot)$ decrease, thus indicates multi-objective Pareto-descent in Definition 3.3.
With $b^{(k)}$ defined above, under $s_{k}$ iterations (number of iterations $s_{k}$ depends on $\theta^{(k)}$), with $b^{(k)}$ fixed, i.e. $b^{(k)}=b^{(k+1)}=\cdots=b^{(k+s_{k}-1)}$, suppose the following constrained optimization Bertsekas (2014) problem starting with $\theta^{(k)}$ has a solution and could be approximated:
$\ell^{(k+s_{k})}(\cdot)=\min_{\mu^{(k+1)},\cdots,\mu^{(k+s_{k})}}\leavevmode\nobreak\ \ell(\cdot)$ (25)
$\quad\mathrm{s.t.}\leavevmode\nobreak$ $R^{(k+s_{k})}(\cdot)\prec b^{(k)}$ (26)
$R^{(j)}(\cdot)\nprec b^{(k)}$ (27)
[…]
$j=k,k+1,\cdots,k+s_{k}-1$ (29)
Then, regardless if the minimization in Equation 25 is attainable or not, as long as we have
$\ell^{(k+m)}(\cdot)$ $<\ell^{(k)}(\cdot),$ (30)
$\exists\leavevmode\nobreak\ m$ $\geq s_{k}$ (31)
, due to Equation 26, Equation 24a, we still achieve multi-objective decrease in Definition 3.3 at step $k+m$ compared to step $k$.
The evolution process of the model parameters in this case can be regarded as a plant with uncertain dynamic $f_{\theta}$ in Equations 14 and 28, where the only control we have at hand is the sequence $\mu^{(k+1)},\cdots,\mu^{(k+s_{k})}$, which provides different loss landscapes at different iterations with different $\theta$ values as depicted in Figure 4. How can we design such $\mu$ or loss landscape sequence to promote Equation 26 to happen?
To simplify the discussion, first consider $\mu,R(\cdot)\in\mathbb{R}_{+}^{d=1}$. Before the constraint in LABEL:{eq:feasibility_r} is satisfied, design $\mu^{(k+1)},\cdots,\mu^{(k+s_{k})}$ to be an increasing sequence Bazaraa et al. (2013), which gives more weight to the corresponding regularization term in $R(\cdot)$, to approximate the constraint optimization in Equation 25 by optimizing Equation 1 for each $\mu$ until Equation 26 is satisfied. However, will the increased weight on the regularization term in $R(\cdot)$ result in deteriorated $\ell(\cdot)$?
Figure 4: Illustration of the concept of multiplier induced loss landscape scheduling we proposed, used as control signal in Figure 1: The lifted landscape $\ell(\cdot)+\mu^{(k+1)}R(\cdot)$ (dotted curve) scheduled at iteration $k+1$, enables the model parameter dynamic to overcome the local minimum $\bar{\theta^{(k)}_{\mu^{(k)}}}$ of the old loss landscape $\ell(\cdot)+\mu^{(k)}R(\cdot)$ (solid curve) scheduled at iteration $k$, as can be imagined via the two balls with different colors rolling along the corresponding loss landscapes. It showcases a scenario corresponding to Definition 3.4: In comparison to $\theta^{(k)}$, $\theta^{(k+1)}$ corresponds to a decreased $\ell(\cdot)+\mu^{(k+1)}R(\cdot)$ value but increased $\ell(\cdot)+\mu^{(k)}R(\cdot)$ value.
One possibility can be, although $f_{\theta}(\mu^{(k+1)},\cdot)$ brings $\theta^{(k)}$ to the next value $\theta^{(k+1)}$ which decreases the penalized loss in Equation 1 with the new $\mu^{(k+1)}$, but the new $\theta^{(k+1)}$ evaluates to a deteriorated penalized loss with respect to the old $\mu^{(k)}$ as shown in Figure 4. To describe this behavior during the increasing process of $\mu$, we have Equation 33 in Definition 3.4. This situation is of particular interest due to the possibility of overcoming local minima of a fixed loss landscape, as indicated in Figure 4.
Definition 3.4.
A reg-Pareto slider $\mathcal{A}(\cdot)$ with respect to function $\ell(\cdot)$ and $R(\cdot)$ is defined to be a set map, s.t.  if $\{\theta^{(k+1)},\mu^{(k+1)}\}\in\mathcal{A}(\mathcal{\ell}(\cdot),R(\cdot);\theta^{(k)},\mu^{(k)})$ then
$\ell(\theta^{(k+1)})+\mu^{(k+1)}R(\theta^{(k+1)})$ $\leq\ell({\theta}^{(k)})+\mu^{(k+1)}R({\theta}^{(k)})$ (32)
$\ell(\theta^{(k+1)})+\mu^{(k)}R(\theta^{(k+1)})$ $\geq\ell({\theta}^{(k)})+\mu^{(k)}R({\theta}^{(k)})$ (33)
$\mu^{(k+1)}$ $\geq\mu^{(k)}$ (34)
In this situation, take $R(\cdot)\in\mathbb{R}^{d=1}$, we have the following conclusion in Corollary 3.1 where we show the increase of $\ell(\cdot)$ as an expense to decrease of $R(\cdot)$ is bounded.
Corollary 3.1.
Let
$\{\theta^{(k+1)},\mu^{(k+1)}\}\in\mathcal{A}(\mathcal{\ell}(\cdot),R(\cdot);\theta^{(k)},\mu^{(k)})$ (35)
from definition 3.4 then,
$R(\theta^{(k+1)})$ $\leq R({\theta}^{(k)})$ (36)
$\ell(\theta^{(k+1)})$ $\geq\ell({\theta}^{(k)})$ (37)
$\ell(\theta^{(k+1)})-\ell({\theta}^{(k)})$ $\leq\mu^{(k+1)}(R({\theta}^{(k)})-R(\theta^{(k+1)}))$ (38)
Proof.
Multiply $-1$ to both sides of Equation 33, add the results to Equation 32, we have
$(\mu^{(k+1)}-\mu^{(k)})R(\theta^{(k+1)})\leq(\mu^{(k+1)}-\mu^{(k)})R({\theta}^{(k)})$ (39)
which implies
$(\mu^{(k+1)}-\mu^{(k)})\left(R(\theta^{(k+1)})-R({\theta}^{(k)})\right)\leq 0$ (40)
Since $\mu^{(k+1)}>\mu^{(k)}$
$R(\theta^{(k+1)})-R({\theta}^{(k)})\leq 0$ (41)
Equation 33 and  32 give
$0\leq\mu^{(k)}(R({\theta}^{(k)})-R(\theta^{(k+1)}))\leq$ (42)
$\ell(\theta^{(k+1)})-\ell({\theta}^{(k)})\leq\mu^{(k+1)}(R({\theta}^{(k)})-R(\theta^{(k+1)}))$ (43)
Remark 3.4.
In the case of $R(\cdot)\in\mathbb{R}^{d>1}$, when all components of $\mu$ increases, Equation 32 and Equation 33 imply at least one component of $R(\cdot)$ will decrease. We defer to Remark 3.6 and Section 3.2.3 for the discussion on when some components of $\mu$ increases while other components decreases.
Remark 3.5.
Definition 3.4 and Corollary 3.1 gives a condition of decreasing $R(\cdot)$ at the expense of increasing $\ell(\cdot)$ (increase bounded though) at iteration $k$. When the gradient component from $\ell(\cdot)$ and $R(\cdot)$ agrees with each other at another iteration, we could expect both objectives to decrease as in Definition 3.3. It can also be the case at a particular iteration, $\ell(\cdot)$ decreases at the expense of $R(\cdot)$ increases.
What is left unclear, however, is after the accumulative effect of several iterations, if $\exists m$ to ensure Equation 30. To answer this, first, Corollary 3.2 gives a bound on the accumulative change of $\ell(\cdot)$ with $m=B$.
Corollary 3.2.
Suppose that for all $k=0,\dots,B-1$
$\ell(\theta^{(k+1)})+\mu^{(k+1)}R(\theta^{(k+1)})$ $\leq\ell(\theta^{(k)})+\mu^{(k+1)}R(\theta^{(k)})$
and $\mu^{(k+1)}>\mu^{(k)}$. Then
$\ell((\theta^{(B)})\leq\ell(\theta^{(0)})+S_{>}+S_{<},$ (44)
where
$S_{>}$ $=\sum_{k\in\mathcal{K}_{>}}\mu^{(k+1)}(R(\theta^{(k)})-R(\theta^{(k+1)})),$
$S_{<}$ $=\sum_{k\in\mathcal{K}_{<}}\mu^{(k)}(R(\theta^{(k)})-R(\theta^{(k+1)})),$
[…]
$\mathcal{K}_{<}$ =$=\{0,\dots,B-1\}\setminus\mathcal{K}_{>}.$
Proof.
For $k\in\mathcal{K}_{>}$ (Equation 33 holds), from Corollary 3.1 we have that
$\ell(\theta^{(k+1)})-\ell(\theta^{(k)})$ $\leq\mu^{(k+1)}(R(\theta^{(k)})-R(\theta^{(k+1)})).$
For $k\in\mathcal{K}_{<}$, it holds that
$\ell(\theta^{(k+1)})+\mu^{(k)}R(\theta^{(k+1)})<\ell(\theta^{(k)})+\mu^{(k)}R(\theta^{(k)}),$
implying
$\ell(\theta^{(k+1)})-\ell(\theta^{(k)})<\mu^{(k)}(R(\theta^{(k)})-R(\theta^{(k+1)})).$
Decomposing
$\ell((\theta^{(B)})-\ell(\theta^{(0)})$
$=\sum_{k=0}^{B-1}[\ell((\theta^{(k+1)})-\ell(\theta^{(k)})]$
$=\Big{(}\sum_{k\in\mathcal{K}_{>}}+\sum_{k\in\mathcal{K}_{<}}\Big{)}[\ell((\theta^{(k+1)})-\ell(\theta^{(k)})]$
yields the conclusion. ∎
Based on Corollary 3.2, we have the following conjecture to ensure Equation 30
Conjecture 3.1.
Further decompose $S_{<}$ from Corollary 3.2 into $S_{<}^{-}$ and $S_{<}^{+}$ via decomposing $\mathcal{K}_{<}$ in Corollary 3.2 into $\mathcal{K}_{<}^{-}$ and $\mathcal{K}_{<}^{+}$. Here $\mathcal{K}_{<}^{-}$ corresponds to the situations of $\ell(\cdot)$ descent with $S_{<}^{-}<0$. And $\mathcal{K}_{<}^{+}$ corresponds to the $\ell(\cdot)$ ascent cases but upperbounded by $S_{<}^{+}=\sum_{k\in\mathcal{K}_{<}^{+}}\mu^{(k)}(R(\theta^{(k)})-R(\theta^{(k+1)}))$. Then $\exists$ sequence $\{\mu^{(k)}\}$ s.t.
$\ell((\theta^{(B)})\leq\ell(\theta^{(0)})+S_{>}+S_{<}^{+}+S_{<}^{-}<0$ (45)
Remark 3.6.
In the case of $d>1$, as long as a loss component of $R(\cdot)^{(k)}$, say $R_{1}^{(k)}(\cdot)$ is not bounded by its corresponding setpoint component $b_{1}^{(k)}(\cdot)$, we can still increase the value of the $\mu$ component $\mu^{(k)}_{1}$ to give more weight to the corresponding loss component, until the loss component in question decreases below the setpoint. However, an increased $\mu$ component $\mu^{(k)}_{1}$ corresponds to more weight on the gradient component corresponding to $R_{1}^{(k)}(\cdot)$, which makes it harder for other components of $R(\cdot)$ and $\ell(\cdot)$ to decrease. Therefore, instead of always increasing each component of $\mu$, if one component of $R$ has overshoot over the setpoint (i.e. constraint switched from being not satisfied to satisfied with respect to a particular loss component), we decrease the $\mu$ value to give the other components of $\mu$ and $R$ more feasible space to adjust themselves.
Remark 3.7.
We could extend the conjecture above to $\exists\mu$ sequence such that after $B$ iterations, Pareto descent in Definition 3.3 could be achieved (equivalently, setpoint should shrink see Remark 3.3). We did observe such a Pareto descent behavior, defined in Definition 3.3 in our experiment for the $d>1$ case of Equation 4, e.g. see Figure 5 and Figure 6.
3.2.3 Output feedback PI-like controller
How exactly should the $\mu$ sequence discussed above look like? With the adaptive law from Equation 23 for setpoint $b$, we design the following controller for $\mu$:
$e^{(k)}$ $=R(\cdot)-b^{(k)}$ (46)
$\delta_{I}^{(k+1)}$ $=(1-\xi_{d})\delta_{I}^{(k)}+\xi_{d}e^{(k)}$ (47)
$\mu^{(k+1)}$ $=\max(\mu^{(k)}\exp^{\max(K_{I}\delta_{I}^{(k+1)},v_{sat})},\mu_{clip})$ (48)
$K_{I}$ $>0$ (49)
$K_{I}$ $\in\mathbb{R}^{d}$ (50)
$k$ $=0,1,\cdots,B$ (51)
In Equation 46, we calculate how far the current output is away from the setpoint $b$, which gets passed through a moving average in Equation 47 with $\xi_{d}$ being the coefficient of moving average in Equation 47 (Rezende and Viola, 2018). $\mu$ is the output of the controller in Equation 48, $K_{I}$ is the control gain for PI (proportional-integration)  Johnson and Moradi (2005) like control, $K_{I}\delta_{I}^{(k+1)}$ is component wise multiplication, $\exp$ and $\max$ are computed component wise, $v_{sat}$ is the exponential shoulder saturation, $\mu_{clip}$ is the upper bound for $\mu$. Note that $\mu_{clip}$ determines the dynamic range of $\mu$ while $v_{sat}$ determines an upper bound about how fast $\mu$ changes.
Remark 3.8.
To avoid arbitrary choice of $K_{I}\in\mathbb{R}^{d}$, we use $\delta^{(0)}=R^{(0)}-b^{(0)}$ in Equation 47 where $b^{(0)}=\rho R^{(0)}$ (a percentage $\rho$). We divide by $\eta v_{sat}$ with $10<\eta<$ 1 to get the value for $K_{I}$ so that the hyperparameter for our algorithm is $\rho,\eta\in\mathbb{R}$ instead of $K_{I}\in\mathbb{R}^{d}$.
The whole process is summarized in Algorithm 1 and Figure 1.
Algorithm 1 Multi-objective Hierarchical Output Feedback Optimization
1:procedure M-Hof-Opt($\mu^{(0)},\theta^{(0)},\rho,\cdot$)
2:     Initialize $\mu^{(0)}$, calculate $b^{(0)}=\rho R^{(0)}$
3:     Compute $K_{I}$ based on Remark 3.8
4:     while budget $B$ not reached do
5:         Update $\mu$ from controller in Equation 48
6:         Update $\theta$ via running Equation 14 with the updated $\mu$ value
7:         Adapt setpoint according to Equation 23.
8:     end while
9:     return $\theta^{(B)}$triangleright
10:end procedure
3.2.4 Multi-objective setpoint based model selection
In general, the feasibility in Equation 26 becomes more difficult to be met each time the setpoint $b$ shrinks. After several setpoint shrinkages, the difficulty of attaining the feasibility can lead to an oscillating behavior of $R(\cdot)$ around the setpoint with uncontrolled amplitude, as shown in Figure 5. To circumvent this behavior, the up-to-event best model is selected at the last setpoint shrinkage.
4 Experiments
Figure 5: Our method drives the different components of $R(\cdot)$ loss terms down at different scales and rates towards the setpoint, which further promotes the setpoint shrinkage. In the Top row, we show the multiplier dynamic as controller output signal in Equation 48 with respect to training epochs. In the Bottom row, we present the tracking behavior of the corresponding regularization loss $R(\cdot)$ in Equation 4 with respect to setpoint $b$ defined in Equation 23. This figure is from training DIVA (Ilse et al., 2020) with our multi-objective hierarchical feedback optimization scheme on the PACS dataset (Li et al., 2017). Left column: For the regularization loss of negative KL-divergence corresponding to multiplier $\beta_{y}$ in  Equation 8 with $R_{\beta_{y}}=-KL(q_{\phi_{y}}(z_{y}|x)||p_{\theta_{y}}(z_{y}|y))$ (blue curve), we see that with the setpoint (red curve) initially set to $\rho=0.99$ in Equation 21 of the initial $R_{\beta_{y}}$ value, in the first few epochs $R_{\beta_{y}}$ overshoots to an even worse value, but our controller pulls it back towards the setpoint in later epochs via increase in $\beta_{y}$ (top figure in the same column) thus adding more weight on $R_{\beta_{y}}(\cdot)$ in the penalized loss. When $R_{\beta_{y}}$ is controlled to be bounded by its setpoint, the corresponding multiplier value $\beta_{y}$ also decays to give space for other objectives to decrease (top figure in the same column). With coordinated $\mu$ dynamic from each component of the controller output, the setpoint shrinks when all $R(\cdot)$ components Pareto dominates the setpoint. After the setpoint shrinks several times, feasibility in Equation 26 becomes harder to maintain and our controller still keeps $R_{\beta_{y}}$ oscillating around the setpoint. Our model selection criteria in Section 3.2.4 selects the model at the last occurrence of the setpoint shrinkage, thus the oscillating behavior at the later epochs can be ignored. Middle column: For the domain classification loss $R_{\gamma_{d}}$ corresponding to multiplier $\gamma_{d}$ in  Equation 10 (blue curve), we observe a similar behavior, but the setpoint shrinkage is more visible. Right column: Although the output component for the loss $R_{\mu_{recon}}$ in Equation 6 oscillates with relatively large amplitude after the last setpoint shrinkage, our model selection criteria at Section 3.2.4 will exclude this oscillating behavior.
Figure 6: Our training scheme adapts the multi-dimensional multiplier $\mu$, steering the optimization process towards a configuration that balances the trade-off between minimizing output $\ell(\cdot)=\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]$ in Equation 5 and output $R_{\gamma_{d}}(\cdot)=\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ in  Equation 10. In this plot, the solid large point corresponds to the initial output. We use gradually changing colors to indicate the training iterations (epochs), indicated by the color bar. For improved visualization, we exclusively utilized the initial 120 epochs and plotted data points solely for every 10 epochs. This figure corresponds to Figure 5.
4.1 Experimental Setting
Figure 7: Our automatic multiplier adjustment scheme ensures robust out-of-domain test accuracy when training DIVA (Ilse et al., 2020) with the modified ResNet50 from (Gulrajani and Lopez-Paz, 2020) on PACS dataset (testing on domain sketch and training on domain photo, art-painting and cartoon). Each panel (i.e., subplot) corresponds to a training scheme. Each box plot inside a panel corresponds to a specific hyperparameter (e.g. multiplier value or controller hyperparameters) combination. Inside each box plot, we repeat the experiment with different random seeds corresponding to the dots scatter. In the first panel, we warm up only $\beta_{y}$, $\beta_{d}$ in Equations 9 and 8 to their ultimate value $1$ while sampling fixed multiplier $\gamma_{d}$ and $\gamma_{y}$ while fixing $\mu_{recon}=1.0$. In the second panel, additionally, we also warm up $\mu_{recon}$ in Equation 6 to $1.0$ and warm up $\gamma_{d}$ in Equation 10 to the sampled value, and keep $\gamma_{y}$ fixed to be the sampled value. In the third panel, we use fixed constant multipliers without warm-up while keep $\beta=1$ and $\mu_{recon}=1$ and only sample $\gamma_{d},\gamma_{y}$ combinations. The last two panels correspond to results utilizing our multi-objective hierarchical feedback optimization training method, and we let $\mu$ include all multipliers except $\gamma_{y}=1$ in Equation 5. For simplicity, when we write $\mu^{(0)}=1.0$ where $\mu^{(0)}\in\mathbb{R}_{+}^{d}$, we mean each component of $\mu^{(0)}$ equals $1.0$.
This section presents our experiments training the domain generalization model DIVA (Ilse et al., 2020) with $6$ loss terms defined in Equation 4 on the widely used domain generalization benchmark dataset PACS (Li et al., 2017) with different training schemes. Given that the sketch domain within the PACS dataset is considered the most inherently challenging, it was chosen as the sole leave-one-out domain to test the performance of domain generalization. We run experiments to answer the following questions:
  • Is our multiplier controller capable of adjusting multidimensional multipliers automatically to drive each component of $R(\cdot)$ at different scales to the setpoint? Does the setpoint shrink, which implies multi-objective descent? We answer these questions with Figure 5 and Figure 6.
  • Will wrong multiplier choices for warmup and fixed multiplier training schemes result in catastrophic effects? Although our method in contrast is multiplier-free, it does need multiplier initial condition $\mu^{(0)}$ and controller hyperparameters. Does varying those have detrimental effects on the predicative performance as well? We answer this question with the benchmark setting in Section 4.1.1 and the results presented in Figure 7.
4.1.1 Benchmark setting
In the benchmark, we sample different combinations of hyperparameters, (including multiplier for baselines and controller hyperparameter for our method) for each method to be compared. For baselines, we sample $\gamma_{d}\in\{1,1001,100001\}$ and $\gamma_{y}\in\{1,1001,100001\}$ value combinations, and set the ultimate values of $\beta_{y},\beta_{d}$ to be $1.0$, following (Ilse et al., 2020). Our method does not need to set ultimate values for $\mu$, so we sample controller hyperparameters, as well as different initial conditions $\mu^{(0)}$ (Each component of the $\mu^{(0)}$ vector set to be the same value).
4.2 Results and Discussion
Regarding the dynamics of how the setpoint $b$ decreases: Initially, Equation 27 holds, in which case we do not update the setpoint, even if some but not all the blue curves (corresponding to $R(\cdot)$ components) are below the red curves (components of the setpoint). After all constraints for each component of $R(\cdot)$ are satisfied as described in Equation 26, the setpoint (red curves) also gets adapted to a new value (which we term shrinkage), as a new goal to be reached by the $R(\cdot)$ loss in Equation 10. Note that the shrinkage of setpoint depends on Pareto dominance in Equation 24a, while non-dominance defined an equivalent relation in Definition 2.2. Our hierarchical output feedback optimization training scheme produced the dynamics of $R(\cdot)$ shown in Figure 5 together with the corresponding setpoints $b$ and multipliers. Note that the multiplier $\beta_{y},\gamma_{d},\mu_{recon}$ and their corresponding loss components operate at different numerical ranges, but our controller still manages to drive the different $R(\cdot)$ loss terms down at different scales and rates.
Figure 6 shows the output portrait of $\ell(\cdot)$ versus a component of $R(\cdot)$ of our training scheme in correspondence to Figure 5. With the multi-objective setpoint based model selection criteria in Section 3.2.4, our method admits a multi-objective descent of the selected model compared to the initial output, such that the uncontrolled behavior at the end of iterations can be safely ignored.
From Figure 7, we can observe that the generalization performance of baselines is highly sensitive to hyperparameter selection. In contrast, our method, as supported by the adaptive behaviors visualized in Figure 5, automatically adjusts the multipliers during the training process, leading to a more robust performance with respect to changes of controller hyperparameters and multiplier initial condition $\mu^{(0)}$. Our method does not need hyperparameter searching of the multiplier, thus saves a fair amount of computational resources.
5 Conclusion
This work addresses the issue of combinatorial choice for multidimensional multipliers weighting many terms in neural network loss function. We proposed a probabilistic graphical model for joint multiplier and model parameter optimization with respect to a multi-objective descent of all loss terms. The estimation of model parameter leads to an automatic adjustment scheme for the multipliers, adopting a hierarchical control scheme, which breaks the multi-objective optimization problem into a series of constraint optimization sub-problems. Each sub-problem is configured with a self-adaptive multi-objective setpoint updated via Pareto dominance. A PI-like multiplier controller drives the loss term to satisfy the setpoint constraint for each sub-problem. Our method operates at the timescale of epoch level during training, thus circumvents the need for exhaustive multiplier search and saves tremendous computational resources compared to methods like Bayesian optimization. Our method demonstrates robust generalization performance in the PACS domain generalization task with Domain Invariant Variational Autoencoding Ilse et al. (2020), compared to other multiplier choice or scheduling schemes which produce very unstable behavior in the training due to the need for a combinatorial choice of multipliers. Our method is multiplier-free since the multiplier is the output of our controller, and it has robust performance with respect to different multiplier initial values. For future work, we are interested in testing the performance of our hierarchical feedback optimization on other domain generalization methods, as well as on other deep learning fields in the form of Equation 1. Theoretically, it is interesting to validate Conjecture 3.1, derive sufficient conditions, incorporate more mathematical analysis into multidimensional case for $R$, and investigate the potential $\mu$ convergence behavior.
6 Contributions And Acknowledgement
XS proposed the idea, developed the theories and algorithms, implemented the code, designed and carried out the experiments, processed experimental data and generated the figures, wrote the manuscript. NC proposed the PI-like multiplier adaptation in Section 3.2.3, helped XS with PI-like controller implementation, held various discussions, proofread the manuscript. AG initiated the visualization code for Figure 5 and Figure 6, refined by XS, proofread and improved the manuscript. YX developed Corollary 3.2 and corresponding remarks with XS, proofread and improved the constraint optimization part of the theory, discussed on control theory part. CF initiated the benchmark code with refinements from XS, proofread Corollary 3.1. ED and FD tested and improved the code for benchmark, proofread the manuscript. FD further improved the tensorboard visualization of the algorithm. LB proofread Corollary 3.1, tested the benchmark code. DS tested the benchmark code, assisted XS in collecting experimental results. CM supervised the project, offered intensive and in depth proofread of the manuscript with major quality improvements.
We sincerely thank Dr. Changxin Liu and Dr. Prof. Jann Rolfes for spending hours on proofreading the manuscript and offering many helpful suggestions to the manuscript. We thank Tomas Chobola for Inkscape support.